OpenStudy (anonymous):

find an equation of the line tangent to \[y=x ^{\sin x}\] at the point x=1

OpenStudy (anonymous):

lny=(sinx)(lnx)

OpenStudy (anonymous):

i'm not even sure where you're starting do i have to use chain rule?

OpenStudy (anonymous):

\[\frac{ 1 }{y }y'=\frac{ sinx }{ x }+lnx(cosx)\]

OpenStudy (anonymous):

implies

OpenStudy (anonymous):

\[y'=y(\frac{ sinx }{ x} +lnx(cosx))\]

OpenStudy (anonymous):

\[y'=x ^{sinx}(\frac{ sinx }{ x }+lnx(cosx))\]

OpenStudy (anonymous):

\[y' at 1 =\sin(1)\]